You hang an object by means of a vertical string. Two strings in turn support this string. String 1 and string 2 make 20o and 60o with the horizontal beam, respectively.

Find the tensions in the strings if the mass of the object is 75.0 kg.


1
Expert's answer
2021-01-13T11:36:21-0500
"T_1\sin{20}=T_2\sin{60}\\75(9.8)=T_1\cos{20}+T_2\cos{60}\\
\\75(9.8)=T_1\cos{20}+T_1\sin{20}\cot{60}\\T_1=646\ N\\T_2=255\ N"